# Beyond Frame-wise Tracking: A Trajectory-based Paradigm for Efficient Point Cloud Tracking

### v1

@@ Abstract
LiDAR-based 3D single object tracking (3D SOT) is a critical task in robotics and autonomous systems. Existing methods typically follow frame-wise motion estimation or a sequence-based paradigm. However, the two-frame methods are efficient but lack long-term temporal context, making them vulnerable in sparse or occluded scenes, while sequence-based methods that process multiple point clouds gain robustness at a significant computational cost. To resolve this dilemma, we propose a novel trajectory-based paradigm and its instantiation, TrajTrack. TrajTrack is a lightweight framework that enhances a base two-frame tracker by implicitly learning motion continuity from historical bounding box trajectories alone—without requiring additional, costly point cloud inputs. It first generates a fast, explicit motion proposal and then uses an implicit motion modeling module to predict the future trajectory, which in turn refines and corrects the initial proposal. Extensive experiments on the large-scale NuScenes benchmark show that TrajTrack achieves new state-of-the-art performance, dramatically improving tracking precision by 4.48% over a strong baseline while running at 56 FPS. Besides, we also demonstrate the strong generalizability of TrajTrack across different base trackers. Code and  video will be available.
@@ I INTRODUCTION
3D single object tracking (3D SOT) is a fundamental task in computer vision with wide applications in autonomous driving and robotics. It aims to accurately track a target across frames using LiDAR or cameras. Compared to camera images, LiDAR point clouds offer notable advantages: robust to light changes and capture rich geometric semantics from environments. These strengths make it well-suited for robust perception in robotics. However, challenges from point cloud inherent sparsity, especially under occlusion, and appearance changes due to motion or blur often lead to incomplete observations, making reliable tracking difficult.
Previous methods typically follow a two-frame paradigm (Fig.1(a)), which can be categorized into appearance-based [1, 2, 3, 4, 5, 6, 7, 8] and motion-based approaches [9, 10, 11, 12, 13, 14, 15]. Similarity-based methods such as SC3D [1], P2B [3], and PTT [6] rely on frame-wise appearance cues between template and search point cloud and perform similarity matching, but they are sensitive to appearance changes and fast motion. In contrast, motion-based methods like $\mathrm{M}^{2}$-Track series [9, 10] and P2P [16] explicitly model inter-frame motion to estimate relative translation. However, by relying solely on two-frame information, they all struggle in sparse or occluded scenes due to limited appearance or motion clues. More critically, they lack considerations of long-term motion continuity, preventing them from forming a predictive motion prior—for example, anticipating a vehicle’s position as it moves through a temporary occlusion—which is essential for robust tracking.
@@ II Related Work
Single Object Tracking on Point Clouds. The field of SOT has evolved significantly, driven by advancements in point cloud processing [19, 20, 21, 22, 23, 24] and the image domain [25, 26, 27]. The pioneering work SC3D [1] applied Siamese networks directly to point clouds, and introduced template matching, but suffered from inefficiency and non-end-to-end training. Addressing these limitations, P2B [3] leveraged a Region Proposal Network (RPN) and VoteNet [28] for real-time proposal generation. 3D-SiamRPN [4] also has a similar pipeline. Follow-up methods like BAT [5] fused box-aware structural features, PTT [6, 7] introduced the transformer, and V2B [29] transformed sparse points into dense bird’s-eye-view maps. Attention mechanisms were also integrated by STNet [30], LTTR [31], and PTTR [8], to propagate target-specific features. Despite progress, appearance-based methods struggled with inefficient geometry in sparse or occlusion scenes. This prompted motion-centric paradigms: $\mathrm{M}^{2}$-Track [9] modeled inter-frame target motion by a ”segmentation and tracking” pipeline. Furthermore, M²Track++ [10] extended the paradigm to semi-supervised applications. To incorporate historical data, DMT [11] predicts the target’s center from past bounding boxes. Recently, P2P[16] considered the context relationship between the background; it inferred the relative motion directly from the cropped point clouds of two consecutive frames. However, these approaches focused on short-term temporal inputs, overlooking long-term historical frames. Recently, TAT[32] addressed this by aggregating historical templates via RNNs but failed to handle low-quality templates. STTracker [17] uses a deformable attention mechanism to aggregate spatio-temporal features from a sequence of BEV feature maps. SeqTrack3D[18] proposed a Sequence-to-Sequence paradigm, considered temporal clues of the target in points and box sequences, and extracted both key geometry and motion information from the point clouds sequence. While recent sequence-based methods improve robustness, they struggle to achieve a better trade-off between accuracy and efficiency due to the high overhead of multi-frame processing and the difficulty of learning consistent motion from complex sequential data.
@@
Sequence Modeling for Motion. Learning motion patterns from temporal sequences is a powerful way to incorporate long-term context. In the broader field of trajectory prediction, VectorNet[33] and LaneRCNN[34] pioneered vectorized/rasterized scene representations with graph neural networks (GNNs) to encode dynamic interactions, enabling an end-to-end manner. GATraj[35] further enhanced interaction modeling through graph attention mechanisms. The TNT[36] and DenseTNT[37] frameworks decomposed prediction into hierarchical target selection and dense trajectory generation, significantly improving long-term rationality. In addition, generative models were used to perform trajectory prediction. For instance, variational autoencoders[38], conditional variational autoencoders (CVAEs) [39], and generative adversarial networks (GANs)[40]. Most of these methods follow a seq2seq structure and predict future trajectories using intermediate features of past trajectories. More recently, Transformer-based architectures have become the dominant approach due to their superior ability to model complex, long-range dependencies in sequential data. For instance, S2TNet[41] demonstrated the power of Transformers for spatio-temporal interactions, while AgentFormer[42] introduced agent-centric attention to effectively model the coherence of an individual’s trajectory over time. CASPFormer[43] integrated causal spatial priors for socially compliant predictions. The proven success of these models in learning rich patterns from sparse positional sequences directly informs our approach. We are inspired from this powerful sequence modeling paradigm for the specific task of 3D SOT, designing a lightweight IMM module to learn motion continuity from an object’s past trajectory.
@@ III Method
Figure 2: Overview of proposal Trajectory-based Paradigm and instantiate TrajTrack. Explicit Motion Proposal uses a two-frame tracking baseline to obtain the local-aware motion proposal. Implicit Trajectory Prediction is used to learn the object’s global-aware motion continuity for trajectory proposal. Finally, Trajectory-guided Proposal Refinement cooperates the local and global-aware motion cues to get the refined 3D BBOX as the final output.
@@ III-B Framework Overview
Our TrajTrack framework is designed to address the challenges of 3D SOT by synergistically combining short-term and long-term motion cues. As illustrated in Fig. 2, it follows a two-stage ”propose-predict-refine” pipeline: (1) Explicit Motion Proposal: First, a two-frame-based explicit motion model rapidly generates an initial, locally-aware tracking proposal for the current frame. (2) Implicit Trajectory Prediction: Second, our novel Implicit Motion Modeling (IMM) module learns the object’s long-term motion continuity by leveraging only the historical bbox prediction, predicting a globally-aware future trajectory. (3) Trajectory-Guided Proposal Refinement: Finally, a post-processing operation from the refinement mechanism uses the trajectory prior to intelligently correct the initial proposal, outputting a more robust tracking result.
@@ III-D Stage 2: Implicit Trajectory Prediction
Motion Continuity Learning. This component’s goal is to learn a rich, latent representation of the object’s motion patterns. We first process the observed past trajectory $\mathbf{X}$ by computing relative displacements $\Delta\mathbf{x}^{t}=\mathbf{x}^{t}-\mathbf{x}^{t-1}$ to extract motion-centric features. These features are fed into a TrajFormer Encoder which models the temporal dependencies and outputs parameters for a prior distribution over the latent variable, $p_{\theta}(\mathbf{Z}|\mathbf{X})=\mathcal{N}(\boldsymbol{\mu}^{p},\mathrm{Diag}(\boldsymbol{\sigma}^{p})^{2})$.
During training, we additionally leverage the ground-truth future trajectory $\mathbf{Y}$. Its features are fused with the past features and processed by the encoder to parameterize a more informed posterior distribution, $q_{\phi}(\mathbf{Z}|\mathbf{Y},\mathbf{X})=\mathcal{N}(\boldsymbol{\mu}^{q},\mathrm{Diag}(\boldsymbol{\sigma}^{q})^{2})$. We enforce consistency between these two distributions by minimizing their Kullback-Leibler (KL) divergence, which regularizes the latent space and encourages the prior to learn expressive, predictive motion patterns.
@@
$\mathbf{f}_{t}=\mathbf{y}_{t}\oplus\mathbf{Z},\quad\mathbf{Z}\sim\begin{cases}q_{\phi}(\mathbf{Z}|\mathbf{X},\mathbf{Y})&\textit{(training)}\\ p_{\theta}(\mathbf{Z})&\textit{(inference)}\end{cases}$ (4)
@@ III-E Trajectory-guided Proposal Refinement
The final post-process of TrajTrack is to dynamiclly fuse the proposals from the explicit and implicit motion models. This stage receives two inputs: the agile but potentially noisy local-aware proposal, $\mathbf{b}_{t}^{\text{local}}$, and the stable, global-aware trajectory proposal, $\mathbf{b}_{t}^{\text{global}}$. To synergize these two, we introduce a confidence-based refinement strategy. We use the Intersection-over-Union (IoU) between the two proposals themselves, IoU($\mathbf{b}_{t}^{\text{local}}$, $\mathbf{b}_{t}^{\text{global}}$), as a metric for the reliability of the short-term, explicit prediction. (1) If the IoU is high (above a threshold $\lambda$), it indicates that both short-term and long-term models are in agreement. In this case, we trust the more precise, local-aware proposal, $\mathbf{b}_{t}^{\text{local}}$, as the final output. (2) If the IoU is low, it signals a potential failure of the explicit model, likely due to sparsity or occlusion. In this scenario, we rely on the more stable, long-term trajectory proposal, $\mathbf{b}_{t}^{\text{global}}$, as a robust fallback. This mechanism allows our tracker to be fast and precise in simple scenarios while leveraging the long-term motion prior to enhance robustness and recover from failures in challenging environments.
III-F Loss Function
Implicit Trajectory Prediction Loss ($\mathcal{L}_{\mathrm{traj}}$). To train our IMM module from Stage 2, we use the standard negative Evidence Lower Bound (ELBO) loss. This loss comprises a reconstruction term, which ensures the predicted trajectory matches the ground truth, and a KL-divergence term, which regularizes the latent space:
TABLE I: Comparison with SOTA methods on the NuScenes dataset. Success and Precision are used for evaluation, Bold and underline denote the best result and the second-best one, respectively.
Method Publish Paradigm Car [64,159] Pedestrian [33,227] Truck [13,587] Trailer [3,352] Bus [2,953] Mean [117,278]
@@ III-F Loss Function
TrajTrack Ours Trajectory 69.28 / 76.72 50.98 / 81.89 68.19 / 69.44 70.70 / 68.02 61.16 / 57.67 63.81 / 76.61
Improvement $\uparrow$ 4.13 / $\uparrow$ 3.82 $\uparrow$ 4.55 / $\uparrow$ 6.81 $\uparrow$ 3.23 / $\uparrow$ 3.48 $\uparrow$ 0.24 / $\uparrow$ 1.16 $\uparrow$ 2.14 / $\uparrow$ 1.11 $\uparrow$ 3.97 / $\uparrow$ 4.48
@@ IV Experiments
Dataset and Metrics. We follow the common setup [9, 49] and conduct experiments on the large-scale nuScenes [50] dataset. Notably, due to the limited data size of the KITTI dataset (only 19 training, 2 validation sequences [7, 46]) makes it challenging to adequately evaluate the methods. In contrast, the nuScenes dataset comprises 700 training and 150 validation sequences, across 40K point cloud frames, allowing for a more comprehensive evaluation. The evaluation metrics is followed the common setup [6, 7, 3] to report Success and Precision based on one pass evaluation (OPE) [51, 52].
Implementation Details. We follow previous works [3, 16, 7], set the previous frame t1t{-}1 is set as the template and the current frame $t$ as the search region. These cropped regions, defined as [(4.8,4.8),(4.8,4.8),(1.5,1.5)][(-4.8,4.8),(-4.8,4.8),(-1.5,1.5)] for cars and [(1.92,1.92),(1.92,1.92),(1.5,1.5)][(-1.92,1.92),(-1.92,1.92),(-1.5,1.5)] for humans along the $(x,y,z)$ axes. Additionally, the data augmentation, such as random horizontal flipping and uniform rotation within [-5,5] is applied to target points and bounding boxes to improve the model’s generalization. All experiments are conducted on NVIDIA RTX 3090 GPUs. The network is trained for 20 epochs with the AdamW optimizer, initialized with a learning rate of 0.0001, with a batch size is 32. Our experiment video is available at https://www.bilibili.com/video/BV1ahYgzmEWP.

### v2

# arXiv:2509.11453v2[cs.CV] 28 Feb 2026

@@ Abstract
LiDAR-based 3D single object tracking (3D SOT) is a critical task in robotics and autonomous systems. Existing methods typically follow frame-wise motion estimation or a sequence-based paradigm. However, the two-frame methods are efficient but lack long-term temporal context, making them vulnerable in sparse or occluded scenes, while sequence-based methods that process multiple point clouds gain robustness at a significant computational cost. To resolve this dilemma, we propose a novel trajectory-based paradigm and its instantiation, TrajTrack. TrajTrack is a lightweight framework that enhances a base two-frame tracker by implicitly learning motion continuity from historical bounding box trajectories alone—without requiring additional, costly point cloud inputs. It first generates a fast, explicit motion proposal and then uses an implicit motion modeling module to predict the future trajectory, which in turn refines and corrects the initial proposal. Extensive experiments on the large-scale NuScenes benchmark show that TrajTrack achieves new state-of-the-art performance, dramatically improving tracking precision by 3.02% over a strong baseline while running at 55 FPS. Besides, we also demonstrate the strong generalizability of TrajTrack across different base trackers. Code is available at https://github.com/FiBonaCci225/TrajTrack.
@@ I INTRODUCTION
Environments perception is important for autonomous driving and mobile robots, as it provides essential contextual information for scene understanding and decision-making [63, 64, 29, 55, 34, 28, 26, 35, 30, 31, 43, 62, 42, 46, 45]. Within this context, 3D single object tracking (3D SOT) has emerged as a key task in computer vision, enabling fine-grained localization and continuous monitoring of individual objects—capabilities that are crucial for real-world applications such as autonomous navigation and robotic interaction [18, 5, 69]. It aims to accurately track a target across frames using LiDAR or cameras. Unlike Multi-object tracking (MOT) [47, 38, 60, 7], which is mostly based on the tracking-by-detection paradigm and highly sensitive to missed detections or identity switching, SOT only requires initial target annotation and can perform end-to-end tracking of the target without a detector, providing higher robustness and making it particularly suitable for robotic scenarios that require continuous tracking of a specified target. Compared to camera images, LiDAR point clouds offer notable advantages: robustness to light changes and the ability to capture rich geometries from environments. These strengths make it well-suited for robust perception in robotics [73, 72]. However, challenges from point clouds’ inherent sparsity, especially under occlusion, and appearance changes due to motion or blur often lead to incomplete observations, making reliable tracking difficult.
Previous methods typically follow a two-frame paradigm (Fig.1(a)), which can be categorized into appearance-based [10, 40, 8, 65, 41, 20, 68] and motion-based approaches [66, 67, 25, 71, 69, 52]. Similarity-based methods such as SC3D [10], P2B [40], and PTT [41] rely on frame-wise appearance cues between template and search point cloud and perform similarity matching, but they are sensitive to appearance changes and fast motion. In contrast, motion-based methods like $\mathrm{M}^{2}$-Track series [66, 67] and P2P [37] explicitly model inter-frame motion to estimate relative translation. However, by relying solely on two-frame information, they all struggle in sparse or occluded scenes due to limited appearance or motion clues. More critically, they lack considerations of long-term motion continuity, preventing them from forming a predictive motion prior—for example, anticipating a vehicle’s position as it moves through a temporary occlusion—which is essential for robust tracking.
@@ II Related Work
Object Tracking on Point Clouds. The field of 3D single-object tracking (SOT) has advanced rapidly with the development of point cloud processing. The pioneering work SC3D [10] applied Siamese networks and template matching directly to point clouds, but suffered from inefficiency and non-end-to-end training. To address this, P2B [40] and 3D-SiamRPN [8] employed RPN and VoteNet [39] for efficient proposal generation. Subsequent methods enhanced feature representation: BAT [65] incorporated box-aware structural features, PTT [41, 20] leveraged transformers, and V2B [15] converted sparse points to dense BEV maps. Attention mechanisms were further integrated by STNet [16], LTTR [4], and PTTR [68] to propagate target-specific features. Appearance-based SOTs remain vulnerable under sparsity and occlusion, motivating motion-centric designs. $\mathrm{M}^{2}$-Track [66] modeled inter-frame motion via a segmentation-and-tracking pipeline, extended to semi-supervised settings by M²Track++ [67]. DMT [51] leveraged historical bounding boxes for center prediction, while P2P [37] inferred relative motion directly from cropped point clouds of consecutive frames. However, these methods primarily rely on short-term temporal inputs, neglecting long-term history. To address this, TAT [23] aggregated historical templates via RNNs but struggled with low-quality inputs. STTracker [6] employed deformable attention to integrate spatio-temporal features from BEV sequences, and SeqTrack3D [27] proposed a Sequence-to-Sequence paradigm to capture both geometry and motion from point cloud sequences. While sequence-based approaches improve robustness, they struggle to balance accuracy and efficiency due to the high cost of multi-frame processing and the complexity of modeling consistent motion.
Similarly, 3D MOT exploits temporal cues but along different lines. Filter-based methods (AB3DMOT [47], SimpleTrack [38]) use Kalman or constant-velocity models implicitly maintain motion through state vectors but struggle with nonlinearity. Query-based tracking-by-attention methods (MUTR3D [60], ADA-Track [7]), propagate object identities across frames via attention yet remain detection-driven and focus on association rather than explicit motion modeling. Crucially, MOT methods reinitialize states from detector outputs each frame; by contrast, SOT bypasses per-frame detection through initial target specification, enabling detector-free, persistent, and end-to-end tracking.
@@
Sequence Modeling for Motion. Learning motion patterns from temporal sequences provides an effective way to model long-term context. In trajectory prediction, VectorNet [9] and LaneRCNN [59] introduced vectorized and rasterized scene representations with graph neural networks to capture dynamic interactions. GATraj [3] further enhanced this through graph attention. TNT [61] and DenseTNT [11] improved long-term reasoning via hierarchical goal selection and dense trajectory generation. Generative approaches such as VAEs [49], CVAEs [17], and GANs [12] learned trajectory distributions in a probabilistic manner. Most of these methods adopt a sequence-to-sequence paradigm, predicting future motion from past trajectory features. Recently, Transformer-based architectures have become dominant due to their ability to model complex temporal dependencies, as demonstrated by S2TNet [2] for spatio-temporal interactions and AgentFormer [58] for agent-centric motion coherence, with CASPFormer [54] introducing causal spatial priors for socially compliant forecasting. The success of these models in learning rich motion patterns from sparse positional sequences directly informs our approach. We are inspired from this powerful sequence modeling paradigm for the specific task of 3D SOT, designing a lightweight IMM module to learn motion continuity from an object’s past trajectory.
@@ III Method
Figure 2: Overview of the proposed Trajectory-Based Paradigm and its instantiation TrajTrack. Explicit Motion Proposal uses a two-frame tracking baseline to obtain the local-aware motion proposal. Implicit Trajectory Prediction is used to learn the object’s global-aware motion continuity for trajectory proposal. Finally, Trajectory-guided Proposal Refinement cooperates the local and global-aware motion cues to get the refined 3D BBOX as the final output.
@@ III-B Framework Overview
Our TrajTrack framework is designed to address the challenges of 3D SOT by synergistically combining short-term and long-term motion cues. As illustrated in Fig. 2, it follows a two-stage ”propose-predict-refine” pipeline: (1) Explicit Motion Proposal: First, a two-frame-based explicit motion model rapidly generates an initial, locally-aware tracking proposal for the current frame. (2) Implicit Trajectory Prediction: Second, our novel Implicit Motion Modeling (IMM) module predicts a globally-aware future trajectory. (3) Trajectory-Guided Proposal Refinement: Finally, a post-processing operation from the refinement mechanism uses the trajectory prior to intelligently correct the initial proposal, outputting a more robust tracking result.
@@ III-D Stage 2: Implicit Trajectory Prediction
Motion Continuity Learning. This component learns a latent representation of motion dynamics. We derive motion-centric features from the observed trajectory $\mathbf{X}$ and encode them with TrajFormer encoder and use a MLP to predict statistical parameters($\mu$, $\sigma$) to obtain prior distribution $p_{\theta}(\mathbf{Z}|\mathbf{X})=\mathcal{N}(\boldsymbol{\mu}^{p},\mathrm{Diag}(\boldsymbol{\sigma}^{p})^{2})$. During training, the TrajFormer decoder integrates the future trajectory $\mathbf{Y}$ as conditional input and fuses it with the observed history to predict the posterior distribution $q_{\phi}(\mathbf{Z}|\mathbf{Y},\mathbf{X})=\mathcal{N}(\boldsymbol{\mu}^{q},\mathrm{Diag}(\boldsymbol{\sigma}^{q})^{2})$. A KL divergence loss enforces consistency between the prior and posterior, enabling the latent variable to capture expressive motion patterns.
@@
$\mathbf{f}_{t}=\mathbf{y}_{t}\oplus\mathbf{Z},\quad\mathbf{Z}\sim\begin{cases}q_{\phi}(\mathbf{Z}|\mathbf{X},\mathbf{Y})&\textit{(training)}\\ p_{\theta}(\mathbf{Z}|\mathbf{X})&\textit{(inference)}\end{cases}$ (4)
@@ III-E Trajectory-guided Proposal Refinement
The final post-process of TrajTrack is to dynamically fuse the proposals from the explicit and implicit motion models. This stage receives two inputs: the agile but potentially noisy local-aware proposal, $\mathbf{b}_{t}^{\text{local}}$, and the stable, global-aware trajectory proposal, $\mathbf{b}_{t}^{\text{global}}$. To synergize these two, we introduce a confidence-based refinement strategy. We use the Intersection-over-Union (IoU) between the two proposals themselves, IoU($\mathbf{b}_{t}^{\text{local}}$, $\mathbf{b}_{t}^{\text{global}}$), as a metric for the reliability of the short-term, explicit prediction. (1) If the IoU is high (above a threshold ${\lambda}_{IoU}$), it indicates that both short-term and long-term models are in agreement. In this case, we trust the more precise, local-aware proposal, $\mathbf{b}_{t}^{\text{local}}$, as the final output. (2) If the IoU is low, it signals a potential failure of the explicit model, likely due to sparsity or occlusion. In this scenario, we rely on the more stable, long-term trajecory proposal, $\mathbf{b}_{t}^{\text{global}}$, as a robust fallback. This mechanism allows our tracker to be fast and precise in simple scenarios while leveraging the long-term motion prior to enhance robustness and recover from failures in challenging environments.
TABLE I: Comparison with SOTA methods on the nuScenes dataset. Success and Precision are used for evaluation, Bold and underline denote the best result and the second-best one respectively. * 64159 indicates the number of instances of cars
Method Publish Paradigm Car [64,159]* Pedestrian [33,227] Truck [13,587] Trailer [3,352] Bus [2,953] Mean [117,278]
@@
TrajTrack Ours Trajectory 68.02 / 75.87 48.32 / 78.78 67.19 / 68.44 70.70 / 68.02 61.16 / 57.67 62.25 / 75.15
Improvement $\uparrow$ 2.87 / $\uparrow$ 2.97 $\uparrow$ 1.89 / $\uparrow$ 3.70 $\uparrow$ 2.23 / $\uparrow$ 2.48 $\uparrow$ 0.24 / $\uparrow$ 1.16 $\uparrow$ 2.14 / $\uparrow$ 1.11 $\uparrow$ 2.41 / $\uparrow$ 3.02
III-F Loss Function
Implicit Trajectory Prediction Loss ($\mathcal{L}_{\mathrm{traj}}$). To train our IMM module from Stage 2, we use the standard negative Evidence Lower Bound (ELBO) loss [21]. This loss comprises a reconstruction term, which ensures the predicted trajectory matches the ground truth, and a KL-divergence term, which regularizes the latent space:
@@ IV Experiments
Dataset and Metrics. We follow the common setup [66, 14] and conduct experiments on the large-scale nuScenes [1] dataset. Notably, due to the limited data size of the KITTI dataset (only 19 training, 2 validation sequences [20, 36]) makes it challenging to adequately evaluate the methods. In contrast, the nuScenes dataset comprises 700 training and 150 validation sequences, across 40K point cloud frames, allowing for a more comprehensive evaluation. The evaluation metrics is followed the common setup [41, 20] to report Success and Precision based on one pass evaluation (OPE) [50, 22].
Implementation Details. We follow previous works [40, 37, 20], set the previous frame t1t{-}1 is set as the template and the current frame $t$ as the search region. These cropped regions, defined as [(4.8,4.8),(4.8,4.8),(1.5,1.5)][(-4.8,4.8),(-4.8,4.8),(-1.5,1.5)] for cars and [(1.92,1.92),(1.92,1.92),(1.5,1.5)][(-1.92,1.92),(-1.92,1.92),(-1.5,1.5)] for humans along the $(x,y,z)$ axes. Additionally, the data augmentation, such as random horizontal flipping and uniform rotation within [-5,5] is applied to target points and bounding boxes to improve the model’s generalization. The parameters of IMM are set to $H=2,T=12,{\lambda}_{IoU}=0.5$. All experiments are conducted on NVIDIA RTX 3090 GPUs. The network is trained for 20 epochs with the AdamW optimizer, initialized with a learning rate of 0.0001, with a batch size is 32. Our experiment video is available at https://www.bilibili.com/video/BV1ahYgzmEWP.